# Hybrid Beamforming with Widely-spaced-array for Multi-user Cross-Near-and-Far-Field Communications

### v1

@@ Abstract
With multi-GHz bandwidth, Terahertz (THz) beamforming has drawn increasing attention in the sixth generation (6G) and beyond communications. Existing beamforming designs mainly focus on a compact antenna array where typical communication occurs in the far-field. However, in dense multi-user scenarios, only relying on far-field angle domain fails to distinguish users at similar angles. Therefore, a multi-user widely-spaced array (MU-WSA) is exploited in this paper, which enlarges the near-field region to introduce the additional distance domain, leading to a new paradigm of cross-near-and-far-field (CNFF) communication. Under this paradigm, the CNFF channel model is investigated, based on which the subarray spacing $d_{s}$ and the number of subarrays $K$ in MU-WSA are optimized to maximize the channel capacity. Then, in sub-connected systems, an alternating optimization (AO) beamforming algorithm is proposed to deal with the special block-diagonal format of the analog precoder. For fully-connected systems, a low-complexity steering-vector reconstruction (SVR)-based algorithm is proposed by constructing specialized steering vectors of MU-WSA. Numerical evaluations show that due to distance domain resolutions, the MU-WSA can improve the SE by over $60$at a power of $20$dBm compared to the compact array. Additionally, the proposed AO algorithm in the SC system can achieve over 80% of the sum (SE) of the FC system while reducing the number of phase shifters by $K^{2}$, thereby lowering power consumption. The SVR algorithm in the FC system can achieve over 95% of the upper bound of SE but takes only 10% of the running time of the singular vector decomposition (SVD)-based algorithms.
@@
Terahertz communications, Multi-user hybrid beamforming, Cross near-and-far-field, Widely-spaced array.
@@ I Introduction
Most existing hybrid beamforming schemes consider far-field transmission where the electromagnetic (EM) wavefront is approximated as a plane such that the planar wave model (PWM) is utilized to characterize the channel [6]. Under this assumption, the fully-connected (FC) and the sub-connected (SC) architectures are widely-explored. In FC where each RF chain connects to all the antennas, near-optimal spectral efficiency (SE) can be obtained [7, 8, 9, 10]. While in SC where each RF chain only connects to a subset of antennas, the power consumption and the hardware complexity can be reduced at the cost of SE [11, 12, 13].
However, only considering the far-field PWM imposes two limitations for THz multi-user communications. On one hand, the number of transmit data streams for each user is upper-bounded by the number of multipath components, which could be very limited at the THz band, e.g., less than 5 [14, 15, 16]. On the other hand, as far-field beamforming can only provide angle domain resolution, it cannot distinguish users located at the same or similar angles [17, 18]. Therefore, far-field beamforming lacks the ability to support ubiquitous connectivity in the 6G vision, e.g. $10^{7}/\text{km}^{2}$ [19]. To tackle these problems, near-field beamforming has been proposed to provide additional spatial degrees of freedom (SDoF) due to the non-negligible spherical wavefront in near-field [20, 21]. Leveraging this property, the authors in [20] proposed different beam focusing schemes for multi-users in the fully digital, hybrid, and DMA architectures. Moreover, the authors in [21] proposed the location division multiple access (LDMA) where focused beams can be generated to point towards a specific location rather than just a particular direction as the far-field beam steering does, thus increasing the spatial multiplexing gain, and enabling distance level resolution for multi-user communication.
@@
Several studies of near-field beamforming assume extremely large number of antenna elements [23, 24, 25] or with uniform linear array (ULA) [26, 27, 28] to increase the near-field region. However, incorporating too many antennas within a limited array space leads to complex hardware configurations, while utilizing ULA may restrict the available spatial dimensions. Instead, a promising beamforming architecture for this problem is to utilize the widely-spaced array (WSA) architecture, which was proposed in our previous work for point-to-point (P2P) communication [14] and also known as sparse array, or line-of-sight (LoS) MIMO [29, 30, 31]. Specifically, the antennas at both the transmitter and the receiver are uniformly divided into $K$ widely-spaced subarrays, i.e., the subarray spacing is larger than half-wavelength. As a result, the effective array aperture can be increased without introducing additional antenna elements, thus enlarging the near-field region while maintaining an acceptable hardware complexity. In multi-user scenarios, however, the available size of user equipment and the number of receiving antennas are often significantly constrained, making it impractical to adopt the WSA. Moreover, a new diagram of cross-field communications needs to be investigated where users residing in both near- and far-field are communicated simultaneously.
@@
To fill this research gap, in this article, we design the hybrid beamforming scheme for multi-user WSA (MU-WSA) where we analyze both the FC and the SC cases. The distinctive contributions of this work are summarized as follows.
@@
  • We analyze the WSA system in multi-user scenarios where the base station is equipped with widely-spaced subarrays such that the near-field region is enlarged without increasing the number of transmit antennas. While at each user, the compact array is adopted to fit the limited size of the user equipment. In this system, a cross-near-and-far-field (CNFF) channel model is established, based on which we analyze the influence of the number of subarrays and the subarray spacing on the sum SE. Specifically, we prove that for a fixed number of subarrays, the channel capacity increases with a larger subarray spacing. Then, based on this result, we propose a search-based algorithm to obtain the optimal design for MU-WSA.
@@
  • We investigate the sub-connected MU-WSA where the analog precoder is shown to have a special block-diagonal form. In this case, we propose a subarray-based alternating optimization (AO) algorithm. Specifically, we decompose the design problem into the analog and digital stages where the block-diagonalization (BD) method is applied in the digital stage to mitigate the interference among users. At the analog stage, we iteratively optimize the precoder matrix of each subarray.
  • We investigate the fully-connected MU-WSA where we propose a low-complexity hybrid beamforming algorithm based on steering vector reconstruction (SVR). Specifically, we apply the far-field steering vector to each subarray in WSA, then reconstruct the vectors based on the CNFF channel model to form a special response vector that is suitable for WSA. This algorithm avoids the utilization of the singular value decomposition (SVD) and can achieve a much lower computational complexity with acceptable performance of the sum SE.
@@
  • We analyze and compare the beam pattern and the sum SE performance of the MU-WSA and the compact array in both 2-D and 3-D cross-field communication systems. We show that the MU-WSA can provide additional distance domain resolution to distinguish the users at the same angle, thus increasing the sum SE by over $60$% at a power of $20$dBm. Furthermore, extensive evaluation has been conducted to numerically investigate the performance of the proposed AO and SVR beamforming algorithms. The results show that the proposed AO algorithm in the SC system can achieve over 80% of the sum (SE) of the FC system while reducing the number of phase shifters by $K^{2}$, thereby lowering power consumption. The SVR algorithm in the FC system can achieve over 95% of the upper bound of SE but takes only 10% of the running time of the singular vector decomposition (SVD)-based algorithms.
@@
The remainder of this paper is organized as follows. In Sec. II, we present the channel model and the system model, based on which we formulate the beamforming design problem. Then, in Sec. III, we analyze the impact of the subarray spacing and the number of subarrays on the spectral efficiency and then design the MU-WSA architecture through a search-based algorithm. Then, in IV, we propose the AO algorithm in the SC system and a low-complexity SVR algorithm in the FC system. Simulation results are shown in Sec. V where we compare the beam pattern of the MU-WSA and compact array architecture in the 2D and 3D scenarios. Then we evaluate the performance of the proposed algorithms. Conclusions are drawn in Sec. VI.
@@
Notations: $\mathbf{A}$ is a matrix, $\mathbf{a}$ is a vector and $a$ is a scalar; $\mathbf{A}(m,n)$ represents the element of $\mathbf{A}$ in row $m$ and column $n$; $\mathbf{I}$ is the identity matrix. $\operatorname{blkdiag}(\mathbf{A}_{1},\cdots,\mathbf{A}_{K})$ is the block diagonal matrix that composed of $\mathbf{A}_{1}$ to $\mathbf{A}_{K}$. $(\cdot)^{T}$ and $(\cdot)^{H}$ represents the transpose and conjugate transpose of a matrix. ||$\cdot$|| ⋅ | and $\|\cdot\|_{F}$ denotes the modulus and the Frobenius norm. $\operatorname{det}(\cdot)$ represents the determinant. $\mathbb{C}^{m\times n}$ is the set of complex-valued matrices of dimension $m\times n$.
@@ II-A Cross Near-and-Far-Field (CNFF) Channel Model
As shown in Fig. 1, we consider a downlink MU-WSA system where the BS serves $U$ users simultaneously. At each user, one compact UPA antenna array is equipped due to the limited size of the user equipment. The base station (BS) is equipped with $K$ widely-spaced compact subarrays such that within each subarray, the antennas are spaced with half-wavelength. Because of the small size of the compact antenna array, the Rayleigh distance between the subarray and each user is considered far less than the communication distance $D$, resulting in the far-field propagation scheme. Therefore, according to the planar-wave assumption, the channel matrix between the $u^{\textrm{th}}$ user and the $k^{\textrm{th}}$ subarray at BS can be represented as [7]
$\mathbf{H}^{uk}=\sum_{p=1}^{N_{p}}|\alpha_{p}|e^{-j\frac{2\pi}{\lambda}D_{p}^{uk}}\mathbf{a}_{rp}^{uk}\left(\phi_{rp}^{uk},\theta_{rp}^{uk}\right)\mathbf{a}_{tp}^{uk}\left(\phi_{tp}^{uk},\theta_{tp}^{uk}\right)^{H},$ (1)
where $N_{p}$ is the number of propagation paths for each user, i.e., $N_{p,u}=N_{p},\forall u$. $|\alpha_{p}|$ denotes the modulus of the complex path gain of the $p^{\textrm{th}}$ multipath, and $D_{p}^{uk}$ denotes the transmission distance between the user and the $k^{\textrm{th}}$ subarray. The vectors $\mathbf{a}_{rp}^{uk}\left(\phi_{rp}^{uk},\theta_{rp}^{uk}\right)^{uk}$ and $\mathbf{a}_{tp}^{uk}\left(\phi_{tp}^{uk},\theta_{tp}^{uk}\right)$ are the receive and transmit array response vectors, where $\phi$ and $\theta$ refer to the azimuth and elevation angle, respectively.
Since the widely-spaced subarrays enlarge the effective array aperture of the whole array, the spherical-wave model should be applied among different subarrays. Specifically, by considering different transmission distances and angles of each subarray, the channel matrix of the $u^{\textrm{th}}$ user can be constructed as a combination of the near-field spherical-wave model (SWM) and the far-field PWM, which is written as [32]
@@
Comparing with conventional PWM where $\operatorname{rank}(\mathbf{H}_{u})=N_{p}$, the upper bound of the rank of $\mathbf{H}_{u}$ increases such that $N_{p}\leq\operatorname{rank}(\mathbf{H}_{u})\leq KN_{p}$, for $K\leq\min\{\frac{N_{t}}{N_{p}},\frac{N_{r}}{N_{p}}\}$.
@@ II-B System Model for MU-WSA
At the BS, the transmitted symbols $\mathbf{s}=\left[\mathbf{s}_{1}^{T},...,\mathbf{s}_{U}^{T}\right]^{T}$ first go through a diagonal power allocation matrix $\mathbf{P}=\operatorname{blkdiag}\left(\mathbf{P}_{1},...,\mathbf{P}_{U}\right)\in\mathbb{C}^{UN_{s}\times UN_{s}}$, where $\mathbf{P}_{u}$ is the $u^{\textrm{th}}$ diagonal power allocation matrix. Let $\|\mathbf{P}\|^{2}_{F}=P_{t}$, where $P_{t}$ is the total power constraint. After power allocation, the symbols are processed by a baseband precoder $\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{L_{t}\times UN_{s}}$, followed by an RF precoder $\mathbf{F}_{\mathrm{RF}}=[\mathbf{F}_{1},\cdots,\mathbf{F}_{K}]$ where $\mathbf{F}_{k}\in\mathbb{C}^{\frac{N_{t}}{K}\times L_{t}}$ represents the analog precoder of each subarray. At the $u^{\textrm{th}}$ user, the received signal is processed by the $u^{\textrm{th}}$ RF combiner $\mathbf{W}_{\mathrm{RF},u}\in\mathbb{C}^{N_{r}\times L_{r}}$, followed by the $u^{\textrm{th}}$ baseband combiner $\mathbf{W}_{\mathrm{BB},u}\in\mathbb{C}^{L_{r}\times N_{s}}$. Therefore the received signal of the $u^{\textrm{th}}$ user is described as
@@
where $\mathbf{H}_{u}\in\mathbb{C}^{N_{r}\times N_{t}}$ denotes the channel matrix of the $u^{\textrm{th}}$ user, and $\mathbf{n}_{u}\in\mathbb{C}^{N_{r}\times 1}$ represents the $u^{\textrm{th}}$ noise vector.

### v2

@@ Abstract
With multi-GHz bandwidth, Terahertz (THz) beamforming has drawn increasing attention in the sixth generation (6G) and beyond communications. Existing beamforming designs mainly focus on a compact antenna array where typical communication occurs in the far-field. However, in dense multi-user scenarios, only relying on far-field angle domain fails to distinguish users at similar angles. Therefore, a multi-user widely-spaced array (MU-WSA) is exploited in this paper, which enlarges the near-field region to introduce the additional distance domain, leading to a new paradigm of cross-near-and-far-field (CNFF) communication. Under this paradigm, the CNFF channel model is investigated, based on which the subarray spacing $d_{s}$ and the number of subarrays $K$ in MU-WSA are optimized to maximize the channel capacity. Then, in sub-connected (SC) systems, an subarray-based alternating optimization (S-AO) beamforming algorithm is proposed to deal with the special block-diagonal format of the analog precoder. For fully-connected (FC) systems, a low-complexity steering vector reconstruction (SVR)-based algorithm is proposed by constructing specialized steering vectors of MU-WSA. Numerical evaluations show that due to distance domain resolutions, the MU-WSA can improve the SE by over $60$at a power of $20$ dBm compared to the compact array. Additionally, the proposed S-AO algorithm in the SC system can achieve over 80% of the sum (SE) of the FC system while reducing the number of phase shifters by $K^{2}$, thereby lowering power consumption. The SVR algorithm in the FC system can achieve over 95% of the upper bound of SE, but it takes only 10% of the running time of the singular value decomposition (SVD)-based algorithms.
@@
Terahertz communications, Multi-user hybrid beamforming, Cross-near-and-far-field, Widely-spaced array.
@@ I Introduction
Most existing hybrid beamforming schemes consider far-field transmission where the electromagnetic (EM) wavefront is approximated as a plane such that the planar wave model (PWM) is utilized to characterize the channel [6]. Under this assumption, the fully-connected (FC) and the sub-connected (SC) architectures are widely-explored. In FC where each RF chain connects to all the antennas, near-optimal spectral efficiency (SE) can be obtained [7, 8, 9, 10]. While in SC where each RF chain only connects to a subset of antennas, the power consumption and the hardware complexity can be reduced at the cost of SE [11, 12, 13, 14, 15].
However, only considering the far-field PWM imposes two limitations for THz multi-user communications. On one hand, the number of transmit data streams for each user is upper-bounded by the number of multipath components, which could be very limited at the THz band, e.g., less than 5 [16, 17, 18]. On the other hand, as far-field beam steering can only provide angle domain resolution, it cannot distinguish users located at the same or similar angles [19, 20]. Consequently, far-field beamforming lacks the ability to support ubiquitous connectivity in the 6G vision, e.g., more than $10^{6}/\text{km}^{2}$ [21]. To tackle these problems, near-field beamforming has been proposed to provide additional spatial degrees of freedom (SDoFs) brought by the non-negligible spherical wavefront in near-field [22, 23]. Leveraging this property, the authors in [22] proposed several multi-users beam focusing schemes for fully digital, hybrid, and DMA architectures, respectively. Moreover, the authors in [23] proposed the location division multiple access (LDMA) where focused beams can be generated to point towards a specific location rather than just a particular direction as the far-field beam steering does, thus increasing the spatial multiplexing gain, and enabling distance level resolution for multi-user communications.
@@
Several studies of near-field beamforming assume extremely large number of antenna elements [25, 26, 27] or with uniform linear array (ULA) [28, 29, 30] to increase the near-field region. However, incorporating too many antennas within a limited array space leads to complex hardware configurations, while utilizing ULA may restrict the available spatial dimensions. Another possible solution to endure a large number of antennas and to increase the spatial multiplexing is the distributed antenna system (DAS) [31, 32]. However, DAS usually requires complicated user association strategy and careful design of coordination between distributed antenna arrays, which can increase the complexity of signal processing [33].
@@
Instead, a promising beamforming architecture for this problem is to utilize the widely-spaced array (WSA) architecture, which is also known as sparse array, or line-of-sight (LoS) MIMO [34, 35, 36]. In our previous work, WSA has been proposed for point-to-point (P2P) communication [16]. Specifically, the antennas at both the transmitter and the receiver are uniformly divided into $K$ widely-spaced subarrays, i.e., the subarray spacing is larger than half-wavelength. As a result, the effective array aperture can be increased without introducing numerous antenna elements, thus enlarging the near-field region while maintaining an acceptable hardware complexity. Moreover, compared to DAS, WSA with one common site can easily support joint beamforming without complex inter-site coordination [37]. Based on WSA, the authors in [37] derived a closed-form expression of the maximum SNR at the BS in uplink systems with one single-antenna receiver. When it comes to multi-user scenarios, the available size of user equipment and the number of receiving antennas are often significantly constrained, making it impractical to adopt the WSA. In this case, the authors in [38] focused on an uplink system with multiple single-antenna users, where the beam focusing pattern was analyzed and a user grouping strategy was designed. However, in downlink multi-user systems, a new diagram of cross-field communications occurs where users residing in both near- and far-field are served simultaneously. In this context, the design of the antenna array, including the number of subarrays and the subarray spacing, as well as the beamforming strategy under different connection types, i.e., fully-connected and sub-connected structures, remain unexplored in the literature.
@@
To fill this research gap, in this paper we design the hybrid beamforming scheme for downlink multi-user WSA (MU-WSA), where we analyze both the FC and the SC cases. The distinctive contributions of this work are summarized as follows.
@@
  • We analyze the WSA system in multi-user scenarios where the base station (BS) is equipped with widely-spaced subarrays such that the near-field region is enlarged without increasing the number of transmit antennas. While at each user, the compact array is adopted to fit the limited size of the user equipment. In this system, a cross-near-and-far-field (CNFF) channel model is established as a simplified yet accurate approximation of the spherical-wave model (SWM). Then, we analyze the influence of the number of subarrays and the subarray spacing on the sum SE. Specifically, we prove that for a fixed number of subarrays, the channel capacity increases with a larger subarray spacing. Based on this result, we derive the closed-form expression of the optimal subarray spacing.
@@
  • We investigate the sub-connected MU-WSA with multi-antenna users. Specifically, we first divide the design problem into the analog and digital stages where the block-diagonalization (BD) method is applied in the digital stage to mitigate the user interference. For the analog precoder, we consider a more general case such that each subarray in MU-WSA can be connected to multiple RF chains rather than only one RF chain as in most SC studies [11, 14, 15, 13, 12]. In this context, we formulate the analog precoder as a special block diagonal matrix where each block is itself a matrix. Then, based on mathematical derivations, we decompose the contribution of each submatrix, and propose a subarray-based alternating optimization (S-AO) algorithm to iteratively optimize the precoder for each subarray.
  • We investigate the fully-connected MU-WSA where we propose a novel low-complexity subarray-based steering vector reconstruction (SVR) method, which extends the conventional far-field steering method to the MU-WSA architecture. By leveraging the CNFF channel model, the new beam steering vector for all subarrays is derived in the MU-WSA, using only the transmission angles and transmission distance of the first subarray as parameters, which enables efficient construction of the overall steering vector for the WSA system. Building upon this novel steering vector, the analog precoder for each user can be easily constructed without complex computations such as singular value decomposition (SVD), and achieve acceptable performance of the sum SE.
@@
  • We analyze and compare the beam pattern and the sum SE performance of the MU-WSA and the compact array in both 2-D and 3-D cross-field communication systems. We show that the MU-WSA can provide additional distance domain resolution to distinguish users at the same angle, thus increasing the sum SE by over $60$% at a power of $20$ dBm. Furthermore, extensive evaluation has been conducted to numerically investigate the performance of the proposed S-AO and SVR beamforming algorithms. The results show that the proposed S-AO algorithm in the SC system can achieve over 80% of the sum (SE) of the FC system while reducing the number of phase shifters by $K^{2}$, thereby lowering power consumption. The SVR algorithm in the FC system can achieve over 95% of the upper bound of SE but takes only 10% of the running time of the SVD-based algorithms.
@@
The remainder of this paper is organized as follows. In Sec. II, we present the channel model and the system model, based on which we formulate the beamforming design problem. Then, in Sec. III, we analyze the impact of the subarray spacing and the number of subarrays on the spectral efficiency, based on which the closed-form expression of the optimal subarray spacing is derived. In Sec. IV, we propose the S-AO algorithm in the SC system and a low-complexity SVR algorithm in the FC system. Simulation results are shown in Sec. V where we first compare the beam pattern of the MU-WSA and compact array architecture in 2-D and 3-D scenarios, and we then evaluate the performance of the proposed algorithms. Conclusions are drawn in Sec. VI.
@@
Notations: $\mathbf{A}$ is a matrix, $\mathbf{a}$ is a vector and $a$ is a scalar; $\mathbf{A}(m,n)$ represents the element of $\mathbf{A}$ in row $m$ and column $n$; $\mathbf{I}$ is the identity matrix. $\operatorname{blkdiag}(\mathbf{A}_{1},\cdots,\mathbf{A}_{K})$ is the block diagonal matrix that composed of $\mathbf{A}_{1}$ to $\mathbf{A}_{K}$. $(\cdot)^{T}$ and $(\cdot)^{H}$ represents the transpose and conjugate transpose of a matrix. ||$\cdot$|| ⋅ | and $\|\cdot\|_{F}$ denotes the modulus and the Frobenius norm. $\operatorname{det}(\cdot)$ represents the determinant. $\lceil a\rceil$ denotes the smallest integer no less than $a$. $\mathbb{C}^{m\times n}$ is the set of complex-valued matrices of dimension $m\times n$.
@@ II-A Cross Near-and-Far-Field (CNFF) Channel Model
As shown in Fig. 1, we consider a downlink MU-WSA system where the BS serves $U$ users simultaneously. At each user, one compact UPA antenna array is equipped due to the limited size of the user equipment. The BS is equipped with $K$ widely-spaced subarrays on the X-Z plane. Within each subarray, the antennas are compactly spaced with half-wavelength. Therefore, due to the small size of the subarray, the Rayleigh distance between the subarray and each user is considered far less than the communication distance, resulting in the far-field propagation scheme. Therefore, according to the planar-wave assumption, the channel matrix between the $u^{\textrm{th}}$ user and the $k^{\textrm{th}}$ subarray at BS can be represented as [7]
$\mathbf{H}^{uk}=\sum_{p=1}^{N_{p}}\alpha_{p}\mathbf{a}_{rp}^{uk}\left(\phi_{rp}^{uk},\theta_{rp}^{uk}\right)\mathbf{a}_{tp}^{uk}\left(\phi_{tp}^{uk},\theta_{tp}^{uk}\right)^{H},$ (1)
where $N_{p}$ is the number of propagation paths for each user, i.e., $N_{p,u}=N_{p},\forall u$. $\alpha_{p}$ denotes the complex path gain of the $p^{\textrm{th}}$ multipath. The vectors $\mathbf{a}_{rp}^{uk}\left(\phi_{rp}^{uk},\theta_{rp}^{uk}\right)^{uk}$ and $\mathbf{a}_{tp}^{uk}\left(\phi_{tp}^{uk},\theta_{tp}^{uk}\right)$ are the receive and transmit array response vectors, where $\phi$ and $\theta$ refer to the azimuth and elevation angle, respectively.
However, since the subarrays are widely-spaced, the effective array aperture of the whole array is increased, which in turn extends the Rayleigh distance and enlarges the near-field region. Consequently, the impact of the spherical wavefront must be taken into account across different subarrays, necessitating the use of the near-field model. In this context, the transmission and reception angles of the subarrays cannot be assumed to be identical, and the transmission distance of antennas in different subarrays must be calculated independently using the near-field SWM. Specifically, the channel between the first antenna of different subarrays and the first antenna at each user can be written as
$\mathbf{H}_{u}^{\text{SWM}}$ $=\sum_{p=1}^{N_{p}}\begin{bmatrix}\alpha_{p}^{u1},&\cdots,&\alpha_{p}^{uK}\end{bmatrix}$ (2a)
$\approx\sum_{p=1}^{N_{p}}|\alpha_{p}|\begin{bmatrix}e^{-j\frac{2\pi}{\lambda}D_{p}^{u1}},&\cdots,&e^{-j\frac{2\pi}{\lambda}D_{p}^{uK}}\end{bmatrix},$ (2b)
where eq. (2b) is obtained since the complex path gain between two antennas $i$ and $l$ can be approximated by $\alpha_{p}^{il}\approx|\alpha_{p}^{11}|e^{-j\frac{2\pi}{\lambda}D_{p}^{il}}$ [39].
As a result, the channel matrix of the $u^{\textrm{th}}$ user is constructed as a combination of the near-field SWM among different subarrays and the far-field PWM within each subarray, which can be written as [39]
@@
Comparing with conventional PWM where $\operatorname{rank}(\mathbf{H}_{u})=N_{p}$, the upper bound of the rank of $\mathbf{H}_{u}$ increases such that $N_{p}\leq\operatorname{rank}(\mathbf{H}_{u})\leq KN_{p}$, for $K\leq\min\{\frac{N_{t}}{N_{p}},\frac{N_{r}}{N_{p}}\}$. While comparing to the SWM model where the channel response for each transmission and reception antenna should be calculated independently, the CNFF model requires fewer parameters, which leads to simpler calculation when designing the architecture in Sec. III and beamforming algorithms in Sec. IV. The overall channel matrix can be written as $\mathbf{H}=[\mathbf{H}_{1}^{T},\cdots,\mathbf{H}_{U}^{T}]^{T}$.
@@ II-B System Model for MU-WSA
At the BS, the transmitted symbols $\mathbf{s}=\left[\mathbf{s}_{1}^{T},...,\mathbf{s}_{U}^{T}\right]^{T}$ first go through a diagonal power allocation matrix $\mathbf{P}=\operatorname{blkdiag}\left(\mathbf{P}_{1},...,\mathbf{P}_{U}\right)\in\mathbb{C}^{UN_{s}\times UN_{s}}$, where $\mathbf{P}_{u}$ is the $u^{\textrm{th}}$ diagonal power allocation matrix. Let $\|\mathbf{P}\|^{2}_{F}=P_{t}$, where $P_{t}$ is the total power constraint. After power allocation, the symbols are processed by a baseband precoder $\mathbf{F}_{\mathrm{BB}}=[\mathbf{F}_{\mathrm{BB},1},\cdots,\mathbf{F}_{\mathrm{BB},U}]\in\mathbb{C}^{L_{t}\times UN_{s}}$ where $\mathbf{F}_{\mathrm{BB},u}$ is the digital precoder for the $u^{\mathrm{th}}$ user, followed by an RF precoder $\mathbf{F}_{\mathrm{RF}}=[\mathbf{F}_{1},\cdots,\mathbf{F}_{K}]$ where $\mathbf{F}_{k}\in\mathbb{C}^{\frac{N_{t}}{K}\times L_{t}}$ represents the analog precoder of each subarray. At the $u^{\textrm{th}}$ user, the received signal is processed by the $u^{\textrm{th}}$ RF combiner $\mathbf{W}_{\mathrm{RF},u}\in\mathbb{C}^{N_{r}\times L_{r}}$, followed by the $u^{\textrm{th}}$ baseband combiner $\mathbf{W}_{\mathrm{BB},u}\in\mathbb{C}^{L_{r}\times N_{s}}$. Therefore the received signal of the $u^{\textrm{th}}$ user is described as
@@
where $\mathbf{H}_{u}\in\mathbb{C}^{N_{r}\times N_{t}}$ denotes the channel matrix of the $u^{\textrm{th}}$ user. $\mathbf{n}_{u}\in\mathbb{C}^{N_{r}\times 1}$ represents the $u^{\textrm{th}}$ noise vector, following the complex Gaussian distribution as $\mathcal{CN}(0,\sigma_{n}^{2})$.